# Curvature fluctuations in baby quantum gravity models

### v1

@@ Abstract
Understanding the microscopic behavior of spacetime is critical for developing a theory of quantum gravity and perhaps solving the cosmological constant problem. In this context, it has been proposed that the quantity of interest is the quantum uncertainty in the Ricci scalar and here we investigate this for a discrete baby quantum gravity model based a single square cell. We find that the averaged Ricci scalar vanishes to leading order but has UV-divergent fluctuations. While this behavior is stable under renormalization, it appears not to be under the introduction of a small cosmological constant.
@@ I Introduction
A compelling case can be made for the necessity of background independence in any theory of quantum gravity [1]. However, as the gravitational background is spacetime itself, this makes quantum gravity fundamentally distinct from other quantum theories for which notions like distance, size, and duration are a priori well defined. And indeed constructing a background-independent quantum theory of spacetime is the first hurdle that faces any quantum gravity program.
Nonetheless, heuristic arguments have been made by Wheeler [2, 3] to suggest that whatever the behavior of spacetime may be at a quantum level, we can estimate the magnitude of curvature fluctuations by assuming quantum behavior at the sub-Planckian level. Indeed, this kind of estimation is of particular interest as it may provide an explanation for the observed small scale of the Hubble constant. Thus, in [4, 5] it was argued that all microscopic gravitational fluctuations gravitate attractively and furthermore that they dominate over the repulsive effects of the standard model vacuum. As the fluctuations are in the spacetime geometry itself, they suggest that there is highly inhomogeneous behavior of the spacetime. The idea is that at an intermediate scale well-above the Planck scale but well-below other relevant scales, we may locally model the spacetime fluctuations as an effective stress-energy source with the local Friedmann equation given by
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(-2\rho_{SM}+(1+3w)\rho_{QG})\,,$
@@
where $\rho_{SM}$ is the local average energy density associated with the vacuum state of the standard model, $\rho_{QG}$ is the local average energy density associated with curvature fluctuations, and $w$ characterizes the effective equation of state of the curvature fluctuations. The local Friedmann equation does not require that the scale factor $a$ is spatially constant as the spatial derivatives cancel out. Moreover, $w>-1/3$ because the gravitational fluctuations are always attractive. Hence, when $\rho_{QG}\gg\rho_{SM}$, solutions to the local Friedmann equation become oscillatory rather than exponential. The idea in [4, 5] is that with each small region of spacetime rapidly oscillating independently according to the local Friedmann equation, the scale factor vanishes when averaged over a region many orders of magnitude larger than a single domain. This gives an argument for, absent any additional higher order parametric resonance effects, a vanishing Hubble constant and thus a vanishing macroscopic cosmological constant.
@@
Such models turn around the cosmological constant problem from trying to construct a model that reduces a large “bare” cosmological constant to the small observed value, to constructing models that deviate slightly from a value that would otherwise be zero. This also dovetails with recent results from the DESI collaboration [6] which suggest that the cosmological constant is better fit by a time-dependent field rather than a constant intrinsic to spacetime itself. If this is the case, then the observed expansion cannot be driven by a cosmological constant term in Einstein’s field equations, and thus we would expect $\Lambda=0$ on the macroscopic scale.
Critical in this discussion is that the energy associated with microscopic curvature fluctuations is assumed to be large in a region of spacetime assumed to be small. To that end, we compute the curvature fluctuations at small scales in the context of discrete quantum gravity models based on small graphs. In principle, we are most interested in the behavior of a small spacetime volume, approximated by a single $3+1$ dimensional “cube” of spacetime $(\mathbb{Z}_{2})^{4}$. However, we believe that we may obtain qualitative information about the behavior of this system by instead investigating a simpler graph model, the $1+1$ dimensional Lorentzian square $(\mathbb{Z}_{2})^{2}$, on which quantum gravity was studied in [7]. Briefly, this is a square graph with the horizontal edges assigned a spacelike or negative metric value or “square length” and the vertical edges a timelike or positive one. Quantum or noncommutative Riemannian geometry can be done on any bi-directed graph, with the arrows spanning the space $\Omega^{1}$ of 1-forms and the metric edge data defining a metric $g\in\Omega^{1}\otimes_{A}\Omega^{1}$, where $A$ is the algebra of functions on the vertices. In the present case, there is a 1-parameter family of torsion-free metric-compatible “quantum Levi-Civita” connections labelled by an angle $\theta$, and a natural Ricci scalar curvature $R$ of this. The resulting Einstein-Hilbert action $S_{g}$ does not, however, depend on the parameter. Details are in [7].
To describe the model, it is convenient to consider the vertices as elements of the group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and to note that $\Omega^{1}$ has a 2-dimensional basis of translation invariant 1-forms $e_{1},e_{2}$. We can then write $g=-ae_{1}\otimes e_{1}+be_{2}\otimes e_{2}$ where the four original “square lengths” on the edges can be viewed as positive metric coefficient functions $a,b$ of a certain form. We then $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-Fourier transform these, which expresses their values at the vertices as
@@
$S_{g}=k_{0}\alpha(k)-\ell_{0}\alpha(\ell),\quad\alpha(k)={8k^{2}\over 1-k^{2}}.$
@@
$Z=4\int_{-1}^{1}\int_{-1}^{1}\int_{\varepsilon}^{\mathcal{L}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k\,{\rm d}\ell\,{\rm d}k_{0}\,{\rm d}\ell_{0}\,k_{0}\ell_{0}\,e^{iS_{g}/G}\,.$
@@
$\langle\mathcal{O}\rangle=\frac{4\int_{-1}^{1}\int_{-1}^{1}\int_{\varepsilon}^{\mathcal{L}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k\,{\rm d}\ell\,{\rm d}k_{0}\,{\rm d}\ell_{0}\,k_{0}\ell_{0}\mathcal{O}\,e^{iS_{g}/G}}{Z}\,.$
Notice that the $\ell,\ell_{0}$ theory is the complex conjugate of the $k,k_{0}$ theory. Hence it is enough to work just with the latter and obtain the other side by complex conjugation. Also note that observables which are odd in $k$ have vanishing correlation functions and if we work with even functions of $k$, we can restrict to $k\in[0,1)$. We can then change variables to $\alpha(k)\in[0,\infty)$ in place of $k=\sqrt{\alpha/(\alpha+8)}$. The $k,k_{0}$ factor of $Z$ can now be computed up to constant factors as
$Z_{1}$ $=\int_{0}^{\infty}\int_{0}^{\infty}{{\rm d}\alpha\over\alpha^{1\over 2}(\alpha+8)^{3\over 2}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k_{0}\,k_{0}e^{\imath k_{0}\alpha/G}$
$=\int_{0}^{\infty}{{\rm d}\alpha\over\alpha^{1\over 2}(\alpha+8)^{3\over 2}}{{\rm d}\over{\rm d}\alpha}\left({e^{\imath\mathcal{L}\alpha/G}-e^{\imath\varepsilon\alpha/G}\over\alpha}\right)$
@@ II The Averaged Ricci Scalar
For the purposes of this article, we are interested in the behavior of the vertex-averaged, $\theta$-averaged Ricci scalar. To that end, we need the following operators:
$R(00)$ $=\frac{2k}{1-k}\frac{1}{\ell_{0}(1+\ell)}$
$R(01)$ $=-\frac{2}{k_{0}(1-k)}\frac{\ell}{1-\ell}$
$R(10)$ $=\frac{2}{k_{0}(1+k)}\frac{\ell}{1+\ell}$
$R(11)$ $=-\frac{2k}{1+k}\frac{1}{\ell_{0}(1-\ell)}\,.$
Now note that because expectation values of odd functions of $k$ or $\ell$ vanish, we may write that
$\left\langle\frac{k}{1-k}\right\rangle$ $=\left\langle\frac{k^{2}}{1-k^{2}}\right\rangle=\frac{1}{8}\langle\alpha\rangle$
$\left\langle\frac{k}{1+k}\right\rangle$ $=-\left\langle\frac{k^{2}}{1-k^{2}}\right\rangle=-\frac{1}{8}\langle\alpha\rangle$
$\left\langle\frac{1}{k_{0}(1-k)}\right\rangle$ $=\left\langle\frac{1}{k_{0}(1-k^{2})}\right\rangle=\left\langle\frac{8+\alpha}{8k_{0}}\right\rangle$
$\left\langle\frac{1}{k_{0}(1+k)}\right\rangle$ $=\left\langle\frac{1}{k_{0}(1-k^{2})}\right\rangle=\left\langle\frac{8+\alpha}{8k_{0}}\right\rangle\,.$
@@
$=\tfrac{i}{32}\operatorname{Im}\left(\langle\alpha\rangle\overline{\left\langle\frac{8+\alpha}{k_{0}}\right\rangle}\right)\,.$
@@
$2048\langle R_{av}^{2}\rangle=2\operatorname{Re}\left(\langle\alpha^{2}\rangle\overline{\langle\tfrac{(\alpha+4)(\alpha+8)}{k_{0}^{2}}\rangle}+4\langle\alpha\rangle\overline{\langle\tfrac{\alpha(8+\alpha)}{k_{0}^{2}}\rangle}\right)\,.$
An interesting observation is that $\langle R_{av}\rangle$ is always imaginary and $\langle R_{av}^{2}\rangle$ is always real. We speculate that this indicates that it is the variance of the Ricci scalar, rather than the Ricci scalar itself, which is a physical observable.
Now, in order to study the divergent behavior of these correlators, we calculate in the large $\mathcal{L}$ and small $\varepsilon$ limit. However, even then, the necessary integrals are non-trivial. In some of these cases, named integrals are obtainable (in the form of Meijer-G functions), but series expansions for these special functions are non-trivial. As such, we used the inverse equation solver RIES [8] to determine the exact form of leading and subleading terms in powers of $\mathcal{L}$ from numerics. When this method was unusable, we instead approximated the highly-oscillatory integrands via Fourier methods described in [9, Section 1]. Finally, when even these methods did not produce convergent integrals, we approximated the integrands using piecewise-stitched Taylor series approximations.
@@
$Z_{1}\approx\tfrac{1}{3}(1+i)\sqrt{\pi G\mathcal{L}^{3}}+\tfrac{1}{16}(1-i)\sqrt{\pi G^{3}\mathcal{L}}\,,$
@@
$\langle k_{0}\rangle$ $\approx\frac{3\mathcal{L}}{5}+\frac{3iG}{160}$
$\langle\alpha\rangle$ $\approx\frac{3iG}{2\mathcal{L}}+\frac{27(1-i)}{35}\sqrt{\frac{G^{3}}{\mathcal{L}^{3}}}$
$\langle\alpha^{2}\rangle$ $\approx\frac{6(i-1)}{\sqrt{\pi}}\sqrt{\frac{G^{3}}{\mathcal{L}^{3}}}+\frac{9G^{2}}{4\mathcal{L}^{2}}$
$\left\langle\frac{8+\alpha}{k_{0}}\right\rangle$ $\approx\frac{24}{\mathcal{L}}-\frac{3(1+i)}{\sqrt{\pi}}\sqrt{\frac{G}{\mathcal{L}^{3}}}$
$\langle\tfrac{(\alpha+4)(\alpha+8)}{k_{0}^{2}}\rangle$ $\approx\frac{24(1+i)}{\sqrt{\pi}}\sqrt{\frac{G}{\varepsilon^{2}\mathcal{L}^{3}}}+\frac{9(i-1)}{2\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}$
$\langle\tfrac{\alpha(\alpha+8)}{k_{0}^{2}}\rangle$ $\approx\frac{24(1+i)}{\sqrt{\pi}}\sqrt{\frac{G}{\varepsilon^{2}\mathcal{L}^{3}}}+\frac{9(i-1)}{2\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}\,.$
@@
$\langle R_{av}\rangle\approx\frac{9iG}{8\mathcal{L}^{2}}-\frac{9i(144\sqrt{\pi}+35)}{2240\sqrt{\pi}}\sqrt{\frac{G^{3}}{\mathcal{L}^{5}}}$
@@
$\langle R_{av}^{2}\rangle\approx\frac{9}{64\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}+\frac{81}{1024\sqrt{\pi}}\sqrt{\frac{G^{5}}{\varepsilon^{2}\mathcal{L}^{7}}}\,.$
@@
$\Delta R_{av}\approx\frac{3}{8\pi^{1/4}}\frac{G^{3/4}}{\varepsilon^{1/2}\mathcal{L}^{5/4}}\,.$
The approximations used here are robust under variation of the interpolations and approximations. We thus have an interesting result: for a fixed IR cutoff, there is a UV divergence in the gravitational fluctuations and a vanishing averaged Ricci scalar.

### v2

# Quantum curvature fluctuations and the cosmological constant in a single plaquette quantum gravity model

@@ Abstract
Understanding the microscopic behavior of spacetime, in particular quantum uncertainty in the Ricci scalar, is critical for developing a theory of quantum gravity and perhaps solving the cosmological constant problem. To test this, we compute this quantity for a simple but exact discrete quantum gravity model based on a single plaquette of spacetime. Our results confirm initial speculations of Wheeler from 1955 in finding a UV divergence in the quantum uncertainty. We further show that this behavior is stable under renormalization, but potentially unstable with the introduction of a cosmological constant, suggesting that a bare cosmological constant is ruled out.
@@ I Introduction
Perhaps the most longstanding problem in fundamental physics is the quantization of gravity. It is hoped that whatever the “correct” theory of quantum gravity might be, such a theory would enable us to solve a variety of open problems in physics. Of interest in this letter is the cosmological constant problem: that the naively-calculated vacuum energy density is 50-122 orders of magnitude larger than the observed dark energy density. Among other things, this disagreement implies either an incredibly fine-tuned cancellation occuring or a subtlty in the theory that has not been taken into account [1].
In recent works by Carlip [2], Wang and Unruh [3], and Wang [4], it is suggested that the problem disappears when one considers the gravitationally attractive nature of microscopic curvature fluctuations. These were heuristically discussed in [5, 6, 7] using Wheeler’s “spacetime foam” with magnitude proposed in these works as being of order $\Delta R\sim\ell_{p}/L^{3}$, where $\ell_{p}$ is the Planck length and $L$ is the size of the region under consideration. With this estimate, ultra-local (i.e. Planckian) quantum gravity fluctuations contribute to an effective attractive energy density which overwhelm the (by comparison) paltry repulsive vacuum energy density associated with particle fields.
The idea in  [2, 3, 4] is that, as the fluctuations are in the spacetime geometry itself, spacetime should be treated as highly inhomogeneous. So, at an intermediate length scale well-above the Planck scale but well-below other relevant scales, we may locally model the spacetime fluctuations as an effective stress-energy source within a local Friedmann equation given by
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(-2\rho_{SM}+(1+3w)\rho_{QG})\,,$ (1)
@@
where $\rho_{SM}$ is the average vacuum energy density of the Standard Model calculated in the typical way and diverging as the UV cutoff goes to infinity (with the usual effective equation of state), $\rho_{QG}$ is the local average energy density associated with curvature fluctuations, and $w$ characterizes the effective equation of state of the curvature fluctuations. Note that the local Friedmann equation does not require that the scale factor $a$ is spatially constant as the spatial derivatives cancel out. Moreover, $w>-1/3$ because the gravitational fluctuations are always attractive. Hence, when $\rho_{QG}\gg\rho_{SM}$, solutions to the local Friedmann equation become oscillatory rather than exponential. The idea is that with each small region of spacetime rapidly oscillating independently according to the local Friedmann equation, the scale factor vanishes when averaged over a region many orders of magnitude larger than a single domain. Rigorously justifying this otherwise-intuitive behavior is quite challenging, see [8, 9, 10, 11] for more details on this so-called “averaging problem”. This gives an argument for, absent any additional higher order parametric resonance effects, a vanishing Hubble constant and thus a vanishing macroscopic cosmological constant.
@@
Such a proposal turns around the cosmological constant problem from trying to construct a model that reduces a large “bare” cosmological constant to the small observed value, to constructing models that deviate slightly from a value that would otherwise be zero. This also dovetails with recent results from the DESI collaboration [12] which suggest that the cosmological constant is better fit by a time-dependent field rather than a constant intrinsic to spacetime itself. If this is the case, then the observed expansion cannot be driven by a cosmological constant term in Einstein’s field equations, and thus we would expect $\Lambda=0$ on the macroscopic scale.
As the implication of such arguments are significant, the primary purpose of this letter is to revisit the previous heuristics about quantum uncertainty in the curvature by means of computations within an exact toy quantum gravity model. The model we use is the quantum Lorentzian geometry of a two-dimensional “plaquette” of spacetime $(\mathbb{Z}_{2})^{2}$ in [13]. In principle, we are most interested in the behavior of a small spacetime volume, modelled as a single $3+1$ dimensional hypercube of spacetime $(\mathbb{Z}_{2})^{4}$. Both dimensional reduction [14, 15] and the use of lattice models [16] are established methods to examine quantum gravity. Moreover, the particular approach we use is noncommutative (or quantum) Riemannian geometry as in [17, 13, 18, 19] applied in the case of a bidirected graph, and is distinct from other approaches such as spin networks [20] or dynamical triangulations [21]. The Gauss-Bonnet theorem making 2D gravity topological does not necessarily apply in noncommutative geometry and indeed the single plaquette model has dynamical modes. Hence, we believe that we may still obtain qualitative information about the behavior of this system by instead working with the more computable $(\mathbb{Z}_{2})^{2}$ case. As a graph geometry, it has no external legs and hence can be considered as without boundary.
This approach replaces the functions on a manifold by the algebra $A$ of functions on the set of vertices, the ‘1-forms’ by a vector space $\Omega^{1}$ with basis $\{\omega_{x\to y}\}$ labelled by the arrows, with exterior derivative ${\rm d}:A\to\Omega^{1}$ the finite differences ${\rm d}f=\sum_{x\to y}(f(y)-f(x))\omega_{x\to y}$. The main difference from regular lattice theory is that we multiply 1-forms and functions by $f\omega_{x\to y}=f(x)\omega_{x\to y}$ and $\omega_{x\to y}f=f(y)\omega_{x\to y}$ to have an exact Leibniz rule, but at the price that functions and 1-forms do not generally commute. A metric is a nondegenerate element $g\in\Omega^{1}\otimes_{A}\Omega^{1}$ and in the graph case amounts to assigning a non-zero real ‘square length’ to every arrow. There is no quantisation or deformation parameter in the model as this is simply an exact finite noncommutative geometry. Instead, the metric values and cutoffs that we will impose on them will provide physical scales. We further require a notion of symmetry on the metric by requiring that the square-lengths are the same for arrows in opposite directions. Thus, for the Lorentzian square[13], the metric data is
with horizontal edges assigned a spacelike or negative value and vertical edges a timelike or positive one. We labelled the vertices as elements of the group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. With respect to the latter, $\Omega^{1}$ has a 2-dimensional basis over $A$ of left-invariant 1-forms $e_{1},e_{2}$ and $g=-ae_{1}\otimes e_{1}+be_{2}\otimes e_{2}$, with the metric data now appearing as certain positive metric coefficients $a,b\in A$ valued as shown. There are also higher differential forms with $e_{1},e_{2}$ forming a Grassmann algebra under the wedge product $\wedge$ and killed by the exterior derivative ${\rm d}$.
Given a metric, one then solves for a torsion-free metric-compatible “quantum Levi-Civita” connection $\nabla:\Omega^{1}\to\Omega^{1}\otimes_{A}\Omega^{1}$, finding in the present case[13] a 1-parameter family labelled by a phase $q=e^{\imath\theta}$. (This phase is a ‘purely quantum’ feature of the geometry since classically the Levi-Civita connection is unique.) Here the left most factor of $\nabla$ can be evaluated against a vector field (in the noncommutative geometry sense) to get a covariant derivative along it). From $\nabla$, there is a canonical Riemann curvature $R_{\nabla}=({\rm d}\otimes{\rm id}-{\rm id}\wedge\nabla)\nabla$ as a 2-form-valued operator on one forms. Viewing the 2-form-values as antisymmetric tensors, one can then trace to get the Ricci scalar curvature $R$. Evaluating the expression for this found in [13] at the four points gives
$R(00)$ $={1\over a_{00}b_{00}}\left(\frac{a_{00}^{2}}{a_{01}}-a_{00}+({1+q^{-1}\over 2})(b_{00}-b_{10})\right),$
$R(01)$ $=-{1\over a_{01}b_{00}}\left(\frac{b_{00}^{2}}{b_{10}}-b_{00}+({1+q\over 2})(a_{00}-a_{01})\right),$
$R(10)$ $=-{1\over a_{00}b_{10}}\left({b_{10}^{2}\over b_{00}}-b_{10}-({1+q\over 2})(a_{00}-a_{01})\right),$
$R(11)$ $={1\over a_{01}b_{10}}\left(\frac{a_{01}^{2}}{a_{00}}-a_{01}-({1+q^{-1}\over 2})(b_{00}-b_{10})\right).$ (2)
For the Einstein-Hilbert action we sum $R$ over the four points but with weight function $ab$ (this is the negative determinant of the matrix of $g$ in our basis). This is chosen as in [13] to eliminate the dependence on $\theta$ and results, as found there, in
$S_{g}=(a_{00}-a_{01})^{2}({1\over a_{00}}+{1\over a_{01}})-(b_{00}-b_{10})^{2}({1\over b_{00}}+{1\over b_{10}}).$ (3)
Next, for calculations, it is convenient to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-Fourier transform the metric coefficients, which amounts to
@@
$S_{g}=k_{0}\alpha(k)-\ell_{0}\alpha(\ell),\quad\alpha(k)={8k^{2}\over 1-k^{2}}.$ (5)
@@
$Z=4\int_{-1}^{1}\int_{-1}^{1}\int_{\varepsilon}^{\mathcal{L}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k\,{\rm d}\ell\,{\rm d}k_{0}\,{\rm d}\ell_{0}\,k_{0}\ell_{0}\,e^{iS_{g}/G}\,.$ (6)
@@
$\langle\mathcal{O}\rangle=\frac{4\int_{-1}^{1}\int_{-1}^{1}\int_{\varepsilon}^{\mathcal{L}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k\,{\rm d}\ell\,{\rm d}k_{0}\,{\rm d}\ell_{0}\,k_{0}\ell_{0}\mathcal{O}\,e^{iS_{g}/G}}{Z}\,.$ (7)
Notice that the $\ell,\ell_{0}$ theory is the complex conjugate of the $k,k_{0}$ theory. Hence it is enough to work just with the latter and obtain the other side by complex conjugation. Also note that $Z$ is real and differentiating with respect to the coupling constant brings down $\imath S_{g}$, hence the expectation value of $S_{g}$ is necessarily imaginary. So we expect and will find that the expectation of the averaged Ricci scalar is similarly imaginary.
Next, we note that observables which are odd in $k$ have vanishing correlation functions and if we work with even functions of $k$, we can restrict to $k\in[0,1)$. We can then change variables to $\alpha(k)\in[0,\infty)$ in place of $k=\sqrt{\alpha/(\alpha+8)}$. The $k,k_{0}$ factor of $Z$ can now be computed up to constant factors as [13]
$Z_{1}$ $=\int_{0}^{\infty}{{\rm d}\alpha\over\alpha^{1\over 2}(\alpha+8)^{3\over 2}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k_{0}\,k_{0}e^{\imath k_{0}\alpha/G}$
$=\int_{0}^{\infty}{{\rm d}\alpha\over\alpha^{1\over 2}(\alpha+8)^{3\over 2}}{{\rm d}\over{\rm d}\alpha}\left({e^{\imath\mathcal{L}\alpha/G}-e^{\imath\varepsilon\alpha/G}\over\alpha}\right)$ (8)
@@ II The Averaged Ricci Scalar
For the purposes of this article, we are interested in the behavior of the vertex-averaged, $\theta$-averaged Ricci scalar to reflect what is seen at much larger scales. Averaging (I) over $\theta$ with uniform measure and converting to field momentum variables, we obtain
$R(00)$ $=\frac{2k}{1-k}\frac{1}{\ell_{0}(1+\ell)},$
$R(01)$ $=-\frac{2}{k_{0}(1-k)}\frac{\ell}{1-\ell},$
$R(10)$ $=\frac{2}{k_{0}(1+k)}\frac{\ell}{1+\ell},$
$R(11)$ $=-\frac{2k}{1+k}\frac{1}{\ell_{0}(1-\ell)}\,.$ (9)
Now note that because expectation values (7) of odd functions of $k$ or $\ell$ vanish, we may write that
$\left\langle\frac{k}{1-k}\right\rangle$ $=\left\langle\frac{k^{2}}{1-k^{2}}\right\rangle=\frac{1}{8}\langle\alpha\rangle,$
$\left\langle\frac{k}{1+k}\right\rangle$ $=-\left\langle\frac{k^{2}}{1-k^{2}}\right\rangle=-\frac{1}{8}\langle\alpha\rangle,$
$\left\langle\frac{1}{k_{0}(1-k)}\right\rangle$ $=\left\langle\frac{1}{k_{0}(1-k^{2})}\right\rangle=\left\langle\frac{8+\alpha}{8k_{0}}\right\rangle,$
$\left\langle\frac{1}{k_{0}(1+k)}\right\rangle$ $=\left\langle\frac{1}{k_{0}(1-k^{2})}\right\rangle=\left\langle\frac{8+\alpha}{8k_{0}}\right\rangle\,.$ (10)
@@
$=\tfrac{i}{32}\operatorname{Im}\left(\langle\alpha\rangle\overline{\left\langle\frac{8+\alpha}{k_{0}}\right\rangle}\right)\,.$ (11)
We use parallel formulae to (II) for $l_{0},l$ but recall that these enter with a complex conjugation. These bring their own $1/8$ factor resulting in the $1/64$.
@@
$\langle R_{av}^{2}\rangle={\operatorname{Re}\over 1024}\left(\langle\alpha^{2}\rangle\overline{\langle\tfrac{(\alpha+4)(\alpha+8)}{k_{0}^{2}}\rangle}+4\langle\alpha\rangle\overline{\langle\tfrac{\alpha(8+\alpha)}{k_{0}^{2}}\rangle}\right)\,.$ (12)
An interesting observation is that while $\langle R_{av}\rangle$ is always imaginary as expected, $\langle R_{av}^{2}\rangle$ is always real. This suggests that it is the variance of the Ricci scalar, rather than the Ricci scalar itself, which is a physical observable.
These correlators still depend on the limits $\mathcal{L}$ and $\mathcal{\varepsilon}$ in (I) and we look at them in a large $\mathcal{L}$ and small $\varepsilon$ limit. However, even then, the necessary integrals are non-trivial. In some of these cases, named integrals are obtainable (in the form of Meijer-G functions), but series expansions for these special functions are non-trivial. As such, we used the inverse equation solver RIES [22] to determine the exact form of leading and subleading terms in powers of $\mathcal{L}$ from numerics. When this method was unusable, we instead approximated the highly-oscillatory integrands via Fourier methods described in [23, Section 1]. Finally, when even these methods did not produce convergent integrals, we approximated the integrands using piecewise-stitched Taylor series approximations.
@@
$Z_{1}\approx\tfrac{1}{3}(1+i)\sqrt{\pi G\mathcal{L}^{3}}+\tfrac{1}{16}(1-i)\sqrt{\pi G^{3}\mathcal{L}}\,,$ (13)
@@
$\langle k_{0}\rangle$ $\approx\frac{3\mathcal{L}}{5}+\frac{3iG}{160},$
$\langle\alpha\rangle$ $\approx\frac{3iG}{2\mathcal{L}}+\frac{27(1-i)}{35}\sqrt{\frac{G^{3}}{\mathcal{L}^{3}}},$
$\langle\alpha^{2}\rangle$ $\approx\frac{6(i-1)}{\sqrt{\pi}}\sqrt{\frac{G^{3}}{\mathcal{L}^{3}}}+\frac{9G^{2}}{4\mathcal{L}^{2}},$
$\left\langle\frac{8+\alpha}{k_{0}}\right\rangle$ $\approx\frac{24}{\mathcal{L}}-\frac{3(1+i)}{\sqrt{\pi}}\sqrt{\frac{G}{\mathcal{L}^{3}}},$
$\langle\tfrac{(\alpha+4)(\alpha+8)}{k_{0}^{2}}\rangle$ $\approx\frac{24(1+i)}{\sqrt{\pi}}\sqrt{\frac{G}{\varepsilon^{2}\mathcal{L}^{3}}}+\frac{9(i-1)}{2\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}},$
$\langle\tfrac{\alpha(\alpha+8)}{k_{0}^{2}}\rangle$ $\approx\frac{24(1+i)}{\sqrt{\pi}}\sqrt{\frac{G}{\varepsilon^{2}\mathcal{L}^{3}}}+\frac{9(i-1)}{2\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}\,.$ (14)
@@
$\langle R_{av}\rangle\approx\frac{9iG}{8\mathcal{L}^{2}}-\frac{9i(144\sqrt{\pi}+35)}{2240\sqrt{\pi}}\sqrt{\frac{G^{3}}{\mathcal{L}^{5}}}$ (15)
@@
$\langle R_{av}^{2}\rangle\approx\frac{9}{64\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}+\frac{81}{1024\sqrt{\pi}}\sqrt{\frac{G^{5}}{\varepsilon^{2}\mathcal{L}^{7}}}\,.$ (16)
@@
$\Delta R_{av}\approx\frac{3}{8\pi^{1/4}}\frac{G^{3/4}}{\varepsilon^{1/2}\mathcal{L}^{5/4}}\,.$ (17)
The approximations used here are robust under variation of the interpolations and approximations. We thus have an interesting result: there is a UV divergence in the gravitational fluctuations for a fixed IR cutoff $\mathcal{L}$ and a vanishing averaged Ricci scalar as $\mathcal{L}\to\infty$.